If you mix 55.6ml of dextrose 50% with 944.4ml of dextrose 5% it would result in 1000 ml of dextrose. Need help finding the percentage n please show work
A. 7.5%
B. 95%
C. 75%
D. 9.5%

Answer by math_tutor2020(3816)


Batch A: 55.6 mL of 50% dextrose
Batch B: 944.4 mL of 5% dextrose

From batch A, there's 55.6*0.50 = 27.8 mL of pure dextrose.
From batch B, there's 944.4*0.05 = 47.22 mL of pure dextrose.

In total there's 27.8 + 47.22 = 75.02 mL of pure dextrose out of 1000 mL of total solution.

Divide the two amounts to find the concentration
75.02/1000 = 0.07502 = 7.502%

It appears that your teacher then rounded that to one decimal place to end up with 7.5%


Answer: Choice A